OpenStudy (anonymous):

Determine if the graphs of the given equations are parallel. y=2x+7 7y+2x=4 what is the slope of the line y=2x+7 what is the slope of the line 7y+2x=4 Are the graphs parallel?

OpenStudy (anonymous):

slope form y = mx + b; m is the slope

OpenStudy (anonymous):

what is the slope of the line 7y+2x=4

OpenStudy (anonymous):

Honestly, I am seriosly lost on this question.. I dont even know where to start

OpenStudy (anonymous):

so for the first one 2 is the slope?

OpenStudy (precal):

if the slopes are the same, the lines are parallel

OpenStudy (anonymous):

ok try solving for y 7y+2x=4

OpenStudy (anonymous):

I feel silly but what would be my first step?

OpenStudy (anonymous):

subtract 2x from both sides

OpenStudy (anonymous):

The slope would be -2?

OpenStudy (anonymous):

for the first?

OpenStudy (anonymous):

nope write the whole thing 1st equation slope is 2

OpenStudy (anonymous):

ok..... so the second would be the same then?

OpenStudy (anonymous):

no, try solving for y, will help u

OpenStudy (anonymous):

oh my ok.. let me work this out.. Just a sec

OpenStudy (anonymous):

It would be 4?

OpenStudy (anonymous):

nope 7y+2x=4 subtract 2x from both sides

OpenStudy (anonymous):

sorry I thought I did.. this is embaressing ... ok just a sec

OpenStudy (anonymous):

For some reason I just keep getting 2

OpenStudy (anonymous):

7y+2x=4 -2x -2x 7y = -2x + 4 is that correct?

OpenStudy (anonymous):

(4-2x)/7

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

correct one extra step y = -2/7x + 4/7 whats the slope

OpenStudy (anonymous):

is it undefined?

OpenStudy (anonymous):

y = mx + b m = -2/7

OpenStudy (anonymous):

I dont see where I went wrong....

OpenStudy (anonymous):